# DiffPose-Animal: A Language-Conditioned Diffusion Framework for Animal Pose Estimation

### v1

@@ 1 Introduction
In the domains of ecological conservation and biological research, animal behavior plays a crucial role in environmental monitoring and ecosystem protection. Estimating animal poses allows for the analysis of natural behaviors, thereby enabling the evaluation of ecological health status [1]. Similarly, in intelligent surveillance systems, the detection of abnormal behaviors, such as recumbency or unconsciousness in target animals, contributes to animal safety monitoring and promotes the advancement of intelligent livestock management in agricultural settings [2]. Furthermore, compared to humans, animals, particularly mammals, generally exhibit more complex three-dimensional structures and movement patterns. Since motion capture and modeling operations are fundamentally based on pose estimation, this technique holds substantial significance for applications including animal virtual reality and image reconstruction [3].
In the field of computer vision, pose estimation is a specific task aimed at locating the joints or predefined keypoints of a target (e.g., the human body) in a given image or video, which can concisely and vividly reflect both the static and dynamic information of the target [4], [5]. It has been widely applied in tasks such as action recognition [6], object tracking [7], and 3D reconstruction [8]. Initially, human pose estimation dominated as the mainstream research focus. With the advancement of hardware computational resources, the release of large-scale datasets, and the rapid development of deep learning methods, research on pose estimation has experienced explosive growth and the variety of detected objects has become increasingly diverse. In particular, animals play a crucial role in ecosystems and their behavioral patterns are closely related to humans, making them a key focus in current pose estimation research [9], [10].
@@ 2.1 Animal pose estimation
Early research in animal pose estimation was primarily constrained by limited datasets, with most studies focusing on addressing data scarcity issues. Some approaches leveraged synthetic data through unsupervised or self-supervised learning paradigms. For instance, CC-SSL [11] employed CAD models to generate synthetic animal data with pseudo-labels, followed by self-supervised training based on these generated annotations. However, such synthetic data exhibited domain shifts in comparison to real-world data and contained inherent noise. To mitigate these limitations, UDA-Animal-Pose [10] introduced a multi-scale domain adaptation module to reduce discrepancies between synthetic and real data, coupled with a coarse-to-fine pseudo-label refinement strategy. Further advancing this direction, ScarceNet [1] proposed a novel semi-supervised learning framework that enhanced pseudo-label reliability, effectively alleviating data scarcity challenges. Other works explored transfer learning and domain generalization strategies for animal pose estimation. For example, AnimalPose [12] developed a cross-domain adaptation method to transfer knowledge from labeled to unlabeled animal species while also leveraging human pose priors, demonstrating empirically validated effectiveness. DeepCut [13], building upon the open-source DeepLabCut [14] framework, established a multi-task training architecture for simultaneous animal keypoint and limb prediction, further showing that ImageNet-pre-trained models could enhance both estimation accuracy and generalization capability. Nevertheless, these methods were typically limited to a narrow subset of species. Recent years have witnessed significant progress with the introduction of large-scale animal pose datasets such as AP-10K [15] and AnimalKingdom [16]. Capitalizing on these developments, AnimalRTPose [17] implements high-frequency feature extraction through its CSPNeXt [18] backbone network, integrates channel attention mechanisms to optimize high/low-frequency feature fusion, and employs spatial pyramid pooling for multi-scale contextual representation. This architecture enables robust feature encoding across varying spatial resolutions, substantially improving adaptability to morphologically diverse species and complex environments.
@@ 2.2 Human pose estimation
Human pose estimation is a fundamental task in computer vision, aiming to localize key anatomical landmarks such as the head, shoulders, elbows, and knees from visual input. Existing human keypoint estimation methods can be categorized into three paradigms: single-stage [19], bottom-up [20], and top-down [21]. In the single-stage paradigm, the poses and bounding boxes of all targets in an image are predicted simultaneously. This paradigm presents significant challenges in prediction accuracy, often failing to meet the requirements of practical applications. Additionally, it typically requires complex models for feature extraction and keypoint decoding. Under the bottom-up method, the model predicts the keypoint locations of all target instances in the input image along with the association information between keypoints. Subsequently, post-processing methods such as grouping are employed to assign keypoints to their corresponding instances based on the association information. However, since the grouping process in this paradigm relies on heuristic methods involving numerous empirical techniques, its performance generally falls short of that of the top-down paradigm. With advancements in foundational research such as object detection [22], the top-down approach has emerged as the dominant paradigm for pose estimation. In this paradigm, a detector first identifies the bounding boxes of target instances, which are then expanded by a certain ratio and cropped. Pose estimation is subsequently performed within these instance-specific bounding boxes. With respect to this, we primarily focus on the top-down paradigm in pose estimation.
@@
Typically, Stacked Hourglass Network [23] adopts a multi-stage symmetrical encoder-decoder architecture with skip connections to capture rich spatial and contextual information across multiple scales. By repeatedly downsampling and upsampling feature maps with skip connections, it enables iterative refinement of pose predictions and effectively models complex body structures. Subsequent methods such as the Cascaded Pyramid Network (CPN) [24] enhance representation learning through multi-scale feature fusion and cascaded refinement, leading to more accurate joint localization. SimpleBaseline [21] streamlines the design by appending deconvolution layers to a ResNet [25] backbone, while HRNet [26] maintains high-resolution representations throughout the network and achieves strong performance through continuous multi-scale feature fusion. To address the limitations of heatmap quantization, SimCC [27] reformulates keypoint regression as two independent 1D classification tasks, enabling sub-pixel level localization accuracy. Transformer-based ViTPose [28] further decouples feature encoding and keypoint prediction by employing a pure Vision Transformer backbone, and supports both heatmap and coordinate classification heads for greater modeling flexibility. In addition to architectural innovations, recent work has focused on improving the quality of supervision. RLE [29] loss models more realistic keypoint distributions as soft targets, mitigating the bias introduced by idealized Gaussian heatmaps. Despite impressive accuracy, top-down methods often suffer from high computational cost due to redundant bounding boxes and large pose variability. To address these issues, DynPose [30] introduces a dynamic inference framework that employs a lightweight router to dispatch pose instances to either a small or large pose estimator based on complexity, achieving a more favorable balance between accuracy and efficiency.
@@ 2.3 Diffusion model
Diffusion models have become a dominant paradigm in generative modeling due to their ability to synthesize high-fidelity data. Denoising Diffusion Probabilistic Models (DDPMs) [31] introduce a Markovian forward process that gradually adds noise to data and a reverse process learned to recover it, yielding impressive results across image and audio domains. However, their slow sampling speed has prompted various improvements. DDIM [32] proposes a non-Markovian deterministic formulation that accelerates inference while preserving generation quality. To improve scalability, Latent Diffusion Models (LDMs) [33] perform diffusion in a compressed latent space, significantly reducing computation and enabling high-resolution image generation. Meanwhile, Diffusion Transformers (DiTs) [34] replace U-Nets [35] with transformer-based backbones, offering superior modeling of long-range dependencies and strong performance on large-scale benchmarks. Beyond architectural innovations, recent developments such as Score-based Generative Models (SGMs) [36] and Consistency Models [37] further enhance efficiency and theoretical grounding. These advances collectively establish diffusion models as a flexible and scalable foundation for diverse generative tasks.
@@ 3.1 Preliminary
Denoising diffusion probabilistic models (DDPMs) [31] are a class of generative models that learn to synthesize data through a gradual denoising process. They consist of two main phases: a forward diffusion process and a reverse denoising process. The framework has demonstrated strong performance in modeling complex data distributions and is increasingly applied to structured prediction tasks, such as image/video synthesis [38], audio processing [39], view synthesis [40] and human pose estimation [19].
@@ 3.2.1 Architecture
Image encoder. To extract multi-scale visual features from raw animal images, we employ a convolutional backbone as the image encoder. Specifically, we utilize HRNet [26] as the backbone due to its effectiveness in preserving spatial resolution and capturing rich semantic information. The extracted hierarchical feature maps serve as the visual input to the multimodal fusion module, where they are integrated with the global textual embeddings generated by the LLM.
Text embedding module. To integrate anatomical priors into the pose estimation process, we design a text embedding module that extracts both global and local semantic features from LLMs. For each animal species, we craft a structured prompt that includes the list of keypoints and requests the LLM (e.g., GPT-4o [41]) to perform two tasks: (1) generate a global description covering the animal’s body structure, locomotion style, and biomechanical characteristics relevant to pose, and (2) provide short semantic descriptions for each keypoint, highlighting their anatomical roles or functions.
The combined textual output is then encoded using a frozen text encoder (e.g., CLIP [42], BERT [43], or all-MiniLM-L6-v2 [44] from SentenceTransformers), producing two types of embeddings: a global feature vector $F_{g}$ representing species-level anatomical context, and a set of local keypoint embeddings $F_{l}\in\mathbb{R}^{N\times d}$ where $N$ is the number of keypoints and $d$ is the embedding dimension. These features are later fused with the visual backbone outputs and heatmap representations to guide the pose generation process with semantically meaningful priors.
@@ 4.1 Datasets
The AP-10K [15] dataset is one of the most diverse large-scale benchmarks for animal pose estimation. It contains 10,015 images across 23 animal species, with each instance annotated with 17 keypoints following a unified anatomical scheme. The dataset provides three different train-validation-test splits. AP-10K covers both domestic and wild animals in varied real-world scenarios, posing challenges such as scale variation, background clutter, and occlusion. It is particularly well-suited for evaluating cross-species generalization performance.
AnimalPose [12] is a benchmark dataset introduced for cross-species animal pose estimation. It contains approximately 5,517 images from five quadruped animal categories: dog, cat, horse, cow, and sheep. Each instance is annotated with 18 keypoints, corresponding to semantically aligned body joints across species. The dataset is relatively small in scale but provides high-quality keypoint annotations, making it suitable for evaluating models in low-data regimes and studying generalization across similar quadrupeds.
Animal Kingdom [16] is a recently introduced, large-scale and fine-grained animal pose dataset comprising over 24,000 images from an extensive taxonomy of 50+ species, spanning mammals, birds, and reptiles. Keypoint annotations are species-specific and vary in number, and the dataset includes both common and rare animals under natural environments. Due to its size, species diversity, and scene complexity, AnimalKingdom serves as a rigorous benchmark for evaluating the scalability and robustness of animal pose estimation methods under long-tailed and open-world settings.
@@ 4.2 Evaluation Metrics
To comprehensively evaluate the performance of animal pose estimation models, we adopt a combination of standard keypoint evaluation protocols commonly used in both human and animal pose estimation literature, including the COCO evaluation protocol [45], Object Keypoint Similarity (OKS), Percentage of Correct Keypoints (PCK), and Area Under Curve (AUC).
@@ 4.4 Performance Comparison Across Multiple Datasets
Figure 2: Visualization results of DiffPose-Animal. From top to bottom: AP10K [15], AnimalPose [12], and AnimalKingdom [16]. These examples showcase the diversity of animal categories, poses, environments, and keypoint configurations across datasets. AP-10K includes various mammals under natural conditions; AnimalPose mainly contains domestic animals such as dogs and horses in relatively constrained scenes; Animal Kingdom spans broader species including birds, amphibians, mammal, reptiles and marine animals, highlighting the challenge of cross-species generalization.
@@
YOLOv8-Pose [46] CSPDarknet-N 45.90 - - - - 57.50
@@
AnimalRTPose-N [17] CSPNeXt-N 45.30 - - - - 58.50
AnimalRTPose-S [17] CSPNeXt-S 50.80 - - - - 63.40
AnimalRTPose-M [17] CSPNeXt-M 56.60 - - - - 69.00
AnimalRTPose-L [17] CSPNeXt-L 62.00 - - - - 71.90
AnimalRTPose-X [17] CSPNeXt-X 65.10 - - - - 74.40
Rtmpose [47] Rtmpose-m 72.20 93.90 78.80 56.90 72.80 -
@@
SimpleBaseline [21] ShuffleNetV1 67.20 - - - - 74.80
YOLOv8-Pose [46] CSPDarknet-N 69.50 - - - - 77.40
AnimalRTPose [17] CSPNeXt-N 72.00 - - - - 80.00
HRNet [26] HRNet-W32 73.20 94.78 81.08 70.75 74.65 76.77
@@
P3-mammals hrnet-w32 [26] HRNet-W32 57.10 -
@@
P3-amphibians hrnet-w32 [26] HRNet-W32 53.58 -
@@
P3-reptiles hrnet-w32 [26] HRNet-W32 51.00 -
@@
P3-birds hrnet-w32 [26] HRNet-W32 76.71 -
@@
P3-fishes hrnet-w32 [26] HRNet-W32 64.06 -
@@
Experimental Results on Animal Kingdom. We evaluate our proposed DiffPose-Animal on the Animal Kingdom dataset under various data configurations and taxonomic groups, and report the results in Table 3. Under the P1 and P2 settings, which follow the original configurations proposed in [16], our method consistently outperforms the HRNet-W32 baseline in terms of PCK@0.05, with slight gains (63.25 vs. 63.23 for P1 and 38.24 vs. 37.41 for P2), and also achieves significantly better AUC scores. Under the P3 setting, we further analyze performance across different animal classes, including mammals, amphibians, reptiles, birds, and fishes. Our method achieves consistent improvements in all subcategories, with the most notable gains in mammals (58.36 vs. 57.10) and amphibians (54.08 vs. 53.58), demonstrating the effectiveness of integrating textual semantic priors in enhancing cross-species generalization. Particularly in challenging categories such as reptiles and fishes, DiffPose-Animal shows superior robustness and accuracy. These results confirm the general applicability and strong performance of our approach across a diverse range of species and configurations.

### v2

@@ 1 Introduction
In the domains of ecological conservation and biological research, animal behavior plays a crucial role in environmental monitoring and ecosystem protection. Estimating animal poses allows for the analysis of natural behaviors, thereby enabling the evaluation of ecological health status [li2023scarcenet]. Similarly, in intelligent surveillance systems, the detection of abnormal behaviors, such as recumbency or unconsciousness in target animals, contributes to animal safety monitoring and promotes the advancement of intelligent livestock management in agricultural settings [wiltshire2023deepwild]. Furthermore, compared to humans, animals, particularly mammals, generally exhibit more complex three-dimensional structures and movement patterns. Since motion capture and modeling operations are fundamentally based on pose estimation, this technique holds substantial significance for applications including animal virtual reality and image reconstruction [xu2023animal3d].
In the field of computer vision, pose estimation is a specific task aimed at locating the joints or predefined keypoints of a target (e.g., the human body) in a given image or video, which can concisely and vividly reflect both the static and dynamic information of the target [10.1145/3603618], [WANG2021103225]. It has been widely applied in tasks such as action recognition [choutas2018potion], object tracking [andriluka2018posetrack], and 3D reconstruction [guler2019holopose]. Initially, human pose estimation dominated as the mainstream research focus. With the advancement of hardware computational resources, the release of large-scale datasets, and the rapid development of deep learning methods, research on pose estimation has experienced explosive growth and the variety of detected objects has become increasingly diverse. In particular, animals play a crucial role in ecosystems and their behavioral patterns are closely related to humans, making them a key focus in current pose estimation research [lauer2022multi], [li2021synthetic].
@@ 2.1 Animal pose estimation
Early research in animal pose estimation was primarily constrained by limited datasets, with most studies focusing on addressing data scarcity issues. Some approaches leveraged synthetic data through unsupervised or self-supervised learning paradigms. For instance, CC-SSL [mu2020learning] employed CAD models to generate synthetic animal data with pseudo-labels, followed by self-supervised training based on these generated annotations. However, such synthetic data exhibited domain shifts in comparison to real-world data and contained inherent noise. To mitigate these limitations, UDA-Animal-Pose [li2021synthetic] introduced a multi-scale domain adaptation module to reduce discrepancies between synthetic and real data, coupled with a coarse-to-fine pseudo-label refinement strategy. Further advancing this direction, ScarceNet [li2023scarcenet] proposed a novel semi-supervised learning framework that enhanced pseudo-label reliability, effectively alleviating data scarcity challenges. Other works explored transfer learning and domain generalization strategies for animal pose estimation. For example, AnimalPose [cao2019cross] developed a cross-domain adaptation method to transfer knowledge from labeled to unlabeled animal species while also leveraging human pose priors, demonstrating empirically validated effectiveness. DeepCut [pishchulin2016deepcut], building upon the open-source DeepLabCut [mathis2018deeplabcut] framework, established a multi-task training architecture for simultaneous animal keypoint and limb prediction, further showing that ImageNet-pre-trained models could enhance both estimation accuracy and generalization capability. Nevertheless, these methods were typically limited to a narrow subset of species. Recent years have witnessed significant progress with the introduction of large-scale animal pose datasets such as AP-10K [yu2021ap] and AnimalKingdom [ng2022animal]. Capitalizing on these developments, AnimalRTPose [WU2025107685] implements high-frequency feature extraction through its CSPNeXt [chen2024cspnext] backbone network, integrates channel attention mechanisms to optimize high/low-frequency feature fusion, and employs spatial pyramid pooling for multi-scale contextual representation. This architecture enables robust feature encoding across varying spatial resolutions, substantially improving adaptability to morphologically diverse species and complex environments.
@@ 2.2 Human pose estimation
Human pose estimation is a fundamental task in computer vision, aiming to localize key anatomical landmarks such as the head, shoulders, elbows, and knees from visual input. Existing human keypoint estimation methods can be categorized into three paradigms: single-stage [tan2024diffusionregpose], bottom-up [cheng2020higherhrnet], and top-down [xiao2018simple]. In the single-stage paradigm, the poses and bounding boxes of all targets in an image are predicted simultaneously. This paradigm presents significant challenges in prediction accuracy, often failing to meet the requirements of practical applications. Additionally, it typically requires complex models for feature extraction and keypoint decoding. Under the bottom-up method, the model predicts the keypoint locations of all target instances in the input image along with the association information between keypoints. Subsequently, post-processing methods such as grouping are employed to assign keypoints to their corresponding instances based on the association information. However, since the grouping process in this paradigm relies on heuristic methods involving numerous empirical techniques, its performance generally falls short of that of the top-down paradigm. With advancements in foundational research such as object detection [chen2023diffusiondet], the top-down approach has emerged as the dominant paradigm for pose estimation. In this paradigm, a detector first identifies the bounding boxes of target instances, which are then expanded by a certain ratio and cropped. Pose estimation is subsequently performed within these instance-specific bounding boxes. With respect to this, we primarily focus on the top-down paradigm in pose estimation.
@@
Typically, Stacked Hourglass Network [newell2016stacked] adopts a multi-stage symmetrical encoder-decoder architecture with skip connections to capture rich spatial and contextual information across multiple scales. By repeatedly downsampling and upsampling feature maps with skip connections, it enables iterative refinement of pose predictions and effectively models complex body structures. Subsequent methods such as the Cascaded Pyramid Network (CPN) [chen2018cascaded] enhance representation learning through multi-scale feature fusion and cascaded refinement, leading to more accurate joint localization. SimpleBaseline [xiao2018simple] streamlines the design by appending deconvolution layers to a ResNet [he2016deep] backbone, while HRNet [sun2019deep] maintains high-resolution representations throughout the network and achieves strong performance through continuous multi-scale feature fusion. To address the limitations of heatmap quantization, SimCC [li2022simcc] reformulates keypoint regression as two independent 1D classification tasks, enabling sub-pixel level localization accuracy. Transformer-based ViTPose [xu2022vitpose] further decouples feature encoding and keypoint prediction by employing a pure Vision Transformer backbone, and supports both heatmap and coordinate classification heads for greater modeling flexibility. In addition to architectural innovations, recent work has focused on improving the quality of supervision. RLE [li2021human] loss models more realistic keypoint distributions as soft targets, mitigating the bias introduced by idealized Gaussian heatmaps. Despite impressive accuracy, top-down methods often suffer from high computational cost due to redundant bounding boxes and large pose variability. To address these issues, DynPose [xu2025dynpose] introduces a dynamic inference framework that employs a lightweight router to dispatch pose instances to either a small or large pose estimator based on complexity, achieving a more favorable balance between accuracy and efficiency.
@@ 2.3 Diffusion model
Diffusion models have become a dominant paradigm in generative modeling due to their ability to synthesize high-fidelity data. Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising] introduce a Markovian forward process that gradually adds noise to data and a reverse process learned to recover it, yielding impressive results across image and audio domains. However, their slow sampling speed has prompted various improvements. DDIM [song2020denoising] proposes a non-Markovian deterministic formulation that accelerates inference while preserving generation quality. To improve scalability, Latent Diffusion Models (LDMs) [rombach2022high] perform diffusion in a compressed latent space, significantly reducing computation and enabling high-resolution image generation. Meanwhile, Diffusion Transformers (DiTs) [peebles2023scalable] replace U-Nets [ronneberger2015u] with transformer-based backbones, offering superior modeling of long-range dependencies and strong performance on large-scale benchmarks. Beyond architectural innovations, recent developments such as Score-based Generative Models (SGMs) [song2020score] and Consistency Models [song2023consistency] further enhance efficiency and theoretical grounding. These advances collectively establish diffusion models as a flexible and scalable foundation for diverse generative tasks.
@@ 3.1 Preliminary
Denoising diffusion probabilistic models (DDPMs) [ho2020denoising] are a class of generative models that learn to synthesize data through a gradual denoising process. They consist of two main phases: a forward diffusion process and a reverse denoising process. The framework has demonstrated strong performance in modeling complex data distributions and is increasingly applied to structured prediction tasks, such as image/video synthesis [ho2022video], audio processing [choi2023diffv2s], view synthesis [chen2023single] and human pose estimation [tan2024diffusionregpose].
@@ 3.2.1 Architecture
Image encoder. To extract multi-scale visual features from raw animal images, we employ a convolutional backbone as the image encoder. Specifically, we utilize HRNet [sun2019deep] as the backbone due to its effectiveness in preserving spatial resolution and capturing rich semantic information. The extracted hierarchical feature maps serve as the visual input to the multimodal fusion module, where they are integrated with the global textual embeddings generated by the LLM.
Text embedding module. To integrate anatomical priors into the pose estimation process, we design a text embedding module that extracts both global and local semantic features from LLMs. For each animal species, we craft a structured prompt that includes the list of keypoints and requests the LLM (e.g., GPT-4o [hurst2024gpt]) to perform two tasks: (1) generate a global description covering the animal’s body structure, locomotion style, and biomechanical characteristics relevant to pose, and (2) provide short semantic descriptions for each keypoint, highlighting their anatomical roles or functions.
The combined textual output is then encoded using a frozen text encoder (e.g., CLIP [radford2021learning], BERT [reimers2019sentence], or all-MiniLM-L6-v2 [yin2024study] from SentenceTransformers), producing two types of embeddings: a global feature vector $F_{g}$ representing species-level anatomical context, and a set of local keypoint embeddings $F_{l}\in\mathbb{R}^{N\times d}$ where $N$ is the number of keypoints and $d$ is the embedding dimension. These features are later fused with the visual backbone outputs and heatmap representations to guide the pose generation process with semantically meaningful priors.
@@ 4.1 Datasets
The AP-10K [yu2021ap] dataset is one of the most diverse large-scale benchmarks for animal pose estimation. It contains 10,015 images across 23 animal species, with each instance annotated with 17 keypoints following a unified anatomical scheme. The dataset provides three different train-validation-test splits. AP-10K covers both domestic and wild animals in varied real-world scenarios, posing challenges such as scale variation, background clutter, and occlusion. It is particularly well-suited for evaluating cross-species generalization performance.
AnimalPose [cao2019cross] is a benchmark dataset introduced for cross-species animal pose estimation. It contains approximately 5,517 images from five quadruped animal categories: dog, cat, horse, cow, and sheep. Each instance is annotated with 18 keypoints, corresponding to semantically aligned body joints across species. The dataset is relatively small in scale but provides high-quality keypoint annotations, making it suitable for evaluating models in low-data regimes and studying generalization across similar quadrupeds.
Animal Kingdom [ng2022animal] is a recently introduced, large-scale and fine-grained animal pose dataset comprising over 24,000 images from an extensive taxonomy of 50+ species, spanning mammals, birds, and reptiles. Keypoint annotations are species-specific and vary in number, and the dataset includes both common and rare animals under natural environments. Due to its size, species diversity, and scene complexity, AnimalKingdom serves as a rigorous benchmark for evaluating the scalability and robustness of animal pose estimation methods under long-tailed and open-world settings.
@@ 4.2 Evaluation Metrics
To comprehensively evaluate the performance of animal pose estimation models, we adopt a combination of standard keypoint evaluation protocols commonly used in both human and animal pose estimation literature, including the COCO evaluation protocol [lin2014microsoft], Object Keypoint Similarity (OKS), Percentage of Correct Keypoints (PCK), and Area Under Curve (AUC).
@@ 4.4 Performance Comparison Across Multiple Datasets
Figure 2: Visualization results of DiffPose-Animal. From top to bottom: AP10K [yu2021ap], AnimalPose [cao2019cross], and AnimalKingdom [ng2022animal]. These examples showcase the diversity of animal categories, poses, environments, and keypoint configurations across datasets. AP-10K includes various mammals under natural conditions; AnimalPose mainly contains domestic animals such as dogs and horses in relatively constrained scenes; Animal Kingdom spans broader species including birds, amphibians, mammal, reptiles and marine animals, highlighting the challenge of cross-species generalization.
@@
YOLOv8-Pose [jocher2023ultralytics] CSPDarknet-N 45.90 - - - - 57.50
@@
AnimalRTPose-N [WU2025107685] CSPNeXt-N 45.30 - - - - 58.50
AnimalRTPose-S [WU2025107685] CSPNeXt-S 50.80 - - - - 63.40
AnimalRTPose-M [WU2025107685] CSPNeXt-M 56.60 - - - - 69.00
AnimalRTPose-L [WU2025107685] CSPNeXt-L 62.00 - - - - 71.90
AnimalRTPose-X [WU2025107685] CSPNeXt-X 65.10 - - - - 74.40
Rtmpose [lu2024rtmo] Rtmpose-m 72.20 93.90 78.80 56.90 72.80 -
@@
SimpleBaseline [xiao2018simple] ShuffleNetV1 67.20 - - - - 74.80
YOLOv8-Pose [jocher2023ultralytics] CSPDarknet-N 69.50 - - - - 77.40
AnimalRTPose [WU2025107685] CSPNeXt-N 72.00 - - - - 80.00
HRNet [sun2019deep] HRNet-W32 73.20 94.78 81.08 70.75 74.65 76.77
@@
P3-mammals hrnet-w32 [sun2019deep] HRNet-W32 57.10 -
@@
P3-amphibians hrnet-w32 [sun2019deep] HRNet-W32 53.58 -
@@
P3-reptiles hrnet-w32 [sun2019deep] HRNet-W32 51.00 -
@@
P3-birds hrnet-w32 [sun2019deep] HRNet-W32 76.71 -
@@
P3-fishes hrnet-w32 [sun2019deep] HRNet-W32 64.06 -
@@
Experimental Results on Animal Kingdom. We evaluate our proposed DiffPose-Animal on the Animal Kingdom dataset under various data configurations and taxonomic groups, and report the results in Table 3. Under the P1 and P2 settings, which follow the original configurations proposed in [ng2022animal], our method consistently outperforms the HRNet-W32 baseline in terms of PCK@0.05, with slight gains (63.25 vs. 63.23 for P1 and 38.24 vs. 37.41 for P2), and also achieves significantly better AUC scores. Under the P3 setting, we further analyze performance across different animal classes, including mammals, amphibians, reptiles, birds, and fishes. Our method achieves consistent improvements in all subcategories, with the most notable gains in mammals (58.36 vs. 57.10) and amphibians (54.08 vs. 53.58), demonstrating the effectiveness of integrating textual semantic priors in enhancing cross-species generalization. Particularly in challenging categories such as reptiles and fishes, DiffPose-Animal shows superior robustness and accuracy. These results confirm the general applicability and strong performance of our approach across a diverse range of species and configurations.